OpenStudy (anonymous):

What is the factored form of the expression? 4x2 – 81y2 A. (2x + 9)(2x – 9) B. (2x + 9y)(2x – 9y) C. (2x + 9y)2 D. (2x – 9y)2

hero (hero):

@monays have you already attempted this one?

OpenStudy (akashdeepdeb):

Use the algebraic properties like: LEARN THESE PROPERTIES! \[(a+b)^2 = a^2 + 2ab + b^2\] \[(a-b)^2 = a^2 - 2ab + b^2\] \[(x+a)(x+b) = x^2 + (a+b)x + ab\] \[a^2-b^2 = (a+b)(a-b)\] Okay? :)